Thank you
Ask Your Teac of to be in a report of a study of ca o S8 224 0 187 143 83 71 165 minimum value for this data set, wil this be the case for every data set? Explain. number of equal values at the The lower quartile is equal to the minimum value for this data set because there area equal to the minimum v
Answer #1

putting above data in ascending order:

s. no x
1 0
2 0
3 0
4 0
5 0
6 0
7 0
8 58
9 71
10 83
11 106
12 110
13 143
14 143
15 165
16 177
17 187
18 187
19 187
20 205
21 212
22 224
23 238
24 238
25 307

lower quartile Q1 =npth value =25*0.25 th value =6.25 th value =0

uppper quartile =npth value =25*0.75 th value =18.75 th value =187 (try 196 if this comes wrong)

interquartile range =187-0 =187 ( try 196 if this comes wrong and revert)

the lower.....because there are a large number of equal values ar the lowest end of the distribution

(please provide options given in fillers)